OpenStudy (egbeach):

Find an equation for the nth term of the arithmetic sequence. -3, -5, -7, -9, ...

OpenStudy (freckles):

arithemtic sequences are well known for their common difference

OpenStudy (freckles):

so first step: find common difference

OpenStudy (haseeb96):

@freckles are you help her or should i help her?

OpenStudy (freckles):

it doesn't matter

OpenStudy (freckles):

go for it

OpenStudy (haseeb96):

okay you carry on

OpenStudy (haseeb96):

i am busy some other areas

OpenStudy (freckles):

ok -3, -5, -7, -9, ... common difference can be found be taking a term and subtracting its previous term

OpenStudy (egbeach):

-2

OpenStudy (freckles):

what is -5-(-3)?

OpenStudy (freckles):

yes that is right

OpenStudy (freckles):

\[a_n=a+d(n-1) \\ \text{ so } d=-2 \\ a_n=a-2(n-1)\] a is suppose to represent the first term

OpenStudy (freckles):

what is your first term

OpenStudy (egbeach):

-3

OpenStudy (freckles):

so replacing a with -3 don't replace a_n with anything a_n thing we are finding \[a_n=-3-2(n-1)\]

OpenStudy (egbeach):

thank you!

OpenStudy (freckles):

and as you can see we get all the terms when pluggin n=1,2,3,... \[a_1=-3-2(1-1) \\ a_1=-3-2(0) \\ a_1=-3 \\ a_2=-3-2(2-1) \\ a_2=-3-2(1) \\ a_2=-3-2 \\ a_2=-5 \] and so on ,,,

OpenStudy (freckles):

we will find find that a_n=-3-2(n-1) that a_3=-7 and a_4=-9

OpenStudy (freckles):

and so on... `